### v2

# Cosmic dynamics and observational constraints in $f(Q)$ gravity with affine equation of state

Abstract
In this study, we explore the cosmological implications of $f(Q)$ gravity with the barotropic fluid satisfying the affine equation of state (EoS). We study the impact of affine EoS on the cosmic evolution in the model and derive the functional form $f(Q)$ in this gravity. We show that the non-metricity scalar may describe the unified scenario in which the universe transits from the decelerated expansion into the accelerated expansion. The model parameters are constrained by the Bayesian analysis based on $\chi^{2}$ minimization technique with the observational data of the cosmic chronometer and supernovae type Ia. The cosmic dynamics based on constrained parameters describe the observable universe evolution. The present day values of the cosmological parameters along with the current age of the universe are compatible with the observations.
Keywords: Dark fluid; EoS; Unified scenario; Observations; Non-metricity
1 Introduction
The validation of accelerated cosmic expansion signifies the beginning of a new era in cosmological study. This phenomena was initially noticed while observing the type Ia supernovae [1, 2]. Subsequently, the universe’s expansion has been confirmed by numerous observations [3, 4, 5]. An unusual fluid/field with negative pressure is a cause of the universe’s late-time accelerated expansion [6, 7]. This kind of fluid/field has been termed as the ‘dark energy’ whose exact nature remains mysterious [8, 6]. In other words, the dark energy, theoretically modeled to generate negative pressure has been believed to be responsible for the universe’s accelerated expansion [9, 10, 11, 12, 13, 14]. The most widely accepted model of the universe is the $\Lambda$ cold dark matter ($\Lambda$CDM) model. In this model, the dark energy is described by the constant vacuum energy density which is visualized as the cosmological constant in the model [6, 7]. This model is consistent with the observations, however, it faces several challenges [15, 6, 7, 13]. The enormous difference between the theoretical and observational values of vacuum energy corresponding to cosmological constant causes the issue of fine tuning [6, 7]. In order to resolve the challenges of the $\Lambda$CDM model, the alternative approaches where the dark energy is a dynamical one is generally explored in cosmological modeling.
The issues related to the cosmic acceleration have been explored in dark energy models of the General relativity (GR) as well as the modified theories of gravity. Finding a modification of GR that allows cosmic acceleration without including any exotic components is an interesting option. One of the most well-known ideas to explain the dark contents of the universe is described by the $f(R)$ gravity [16]. This theory is a modification of General Relativity, where $R$ is the Ricci Scalar. Different theories of gravity have been explored to describe issues of dark energy and dark matter, see [17, 18, 19, 20, 10, 11, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and references therein.
In the context of modified gravity theories, Jimenez et al. [33] proposed a gravity theory based on non-metricity and is termed as $f(Q)$ gravity. This theory is based on an approach which considers a variation of the symmetric teleparallel equivalence to General Relativity, here $Q$ is a non-metricity scalar. The studies [13, 34], suggest that the $f(Q)$ theory is one of the intriguing alternative gravity interpretations for explaining cosmological evolution. Lazkoz et al. [35] investigated the accelerated expansion of the universe in context of its consistency with observational data. The exact cosmic solutions that are both anisotropic and isotropic were studied by Esposito et al. [36]. Harko et al. [37] developed a class of $f(Q)$ theories where $Q$ is non-minimally associated with the matter Lagrangian. In connection with the $f(Q)$ gravity theory, numerous further studies [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] suggest that the theory may have interesting cosmological and astrophysical implications, for review see [14].
In this paper, we study the implications of $f(Q)$ gravity with the dark fluid satisfying affine equation of state. We search for an answer whether the $f(Q)$ gravity may have fluid description or not? In particular, for an affine equation of state of dark fluid, we derive the functional formulation of $f(Q)$ gravity. The issue of observational compatibility of model may imply for the observational constraints on model parameters. The dark fluid have been studied in the isotropic and anisotropic spacetime with the General relativity framework [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. The affine equation of state describes the hydro-dynamically stable as well as unstable fluid propagation. The speed of sound remains constant for the fluid. The stable dark energy evolution may be visualized with the fluid satisfying affine equation of state given by $p_{d}=\alpha\rho_{d}-\rho_{0}$, where $\alpha$ and $\rho_{0}$ are constants. The classical stability criterion constrains $0\leq\alpha\leq 1$. In a non-conservative theory, this kind of EoS may naturally arise near the bouncing instant of the bouncing universe [70, 72]. The low energy limits of this kind of EoS may have the $\Lambda$CDM limit [64]. Based on these peculiar properties, we aim to study the cosmological implications of this dark fluid in $f(Q)$ gravity to describe the fluid interpretation of this gravity.
This paper is organized in the following way: An overview of the mathematical aspects of $f(Q)$ gravity in a flat FLRW background has been given in section (2). In Section (3), we obtain the cosmological solutions of the affine EoS model with $f(Q)$ gravity. In Section (4), the Bayesian statistical tools are used to investigate the compatibility of cosmological solutions with the cosmic chronometer (CC) and supernovae type Ia (Pantheon) datasets. The cosmological dynamics based on the parameters such as the deceleration parameter, pressure, energy density, cosmographic parameter and the age of the universe are studied in section (5). The conclusion with summary of results are given in Section (6).
2 An overview of $f(Q)$ gravity
The $f(Q)$ gravity is based on the non-metricity scalar $Q$. The $f(Q)$ gravity is characterized by the following action [33]
${}S=\int\sqrt{-g}\,d^{4}x\left[-\frac{1}{2}f(Q)+\mathcal{L}_{m}\right].$ (1)
Here, $f(Q)$ represents an arbitrary function of the non-metricity scalar $Q$. The symbol $g$ denotes metric tensor’s determinant $g_{ab}$ and the Lagrangian density for matter is symbolized by $\mathcal{L}_{m}$. The quantity $Q$ can be characterized by two distinct traces as
${}Q_{\sigma}=Q_{\sigma}^{\phantom{\sigma}a}a_{\phantom{a}},\quad\tilde{Q}_{\sigma}=Q^{a\phantom{\mu}}_{\phantom{\mu}\sigma a}.$ (2)
Furthermore, the non-metricity scalar $(Q)$ may be expressed by the contraction of above tensors and may be defined as:
${}Q=-Q_{\sigma ab}P^{\sigma ab}.$ (3)
Additionally, the super-potential tensor (which is a conjugate of non-metricity) can be introduced as:
${}4P^{\sigma}_{\phantom{\sigma}ab}=2Q_{(a\phantom{\sigma}b)}^{\phantom{\mu}\sigma}-\tilde{Q}^{\sigma}g_{ab}-Q^{\sigma}_{\phantom{\sigma}ab}-Q^{\sigma}g_{ab}-\delta^{\sigma}_{(a}Q_{b)}^{\phantom{\nu)}}.$ (4)
The field equation may obtained by varying the action (1) with respect to the metric and are expressed as [33]
${}\frac{2}{\sqrt{-g}}\nabla_{\sigma}\left(\sqrt{-g}f_{Q}P^{\sigma}_{ab}\right)+\frac{1}{2}g_{ab}f+f_{Q}\left(P_{a\sigma\beta}Q^{\sigma\beta}_{\phantom{\sigma\beta}b}-2Q_{\sigma\beta a}P^{\sigma\beta}_{\phantom{\sigma\beta}b}\right)=T_{ab},$ (5)
where, $f_{Q}\equiv\frac{df(Q)}{dQ}$ and, the energy-momentum tensor may be written as $T_{ab}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}\mathcal{L}_{m})}{\delta g^{ab}}$. We are assuming the unit system where $18\pi G=c=$ 1.
The energy-momentum tensor $T_{ab}$ might also be written for perfect fluid as $T_{ab}=pg_{ab}+(p+\rho)u_{a}u_{b}$. The perfect fluid’s isotropic pressure and energy density are denoted as $p$ and $\rho$ respectively. The fluid’s $4$-velocity $u_{a}$ satisfies the normalization condition $u_{a}u^{a}=-1$.
Here, we proceed with the framework of spatially flat FLRW spacetime, defined by
${}ds^{2}=a^{2}(t)\left(dx^{2}+dy^{2}+dz^{2}\right)-dt^{2}.$ (6)
Under these conditions, the non-metricity scalar may be described as $Q=6H^{2}$ [33], where $H\equiv\frac{\dot{a}}{a}$ symbolizes the Hubble parameter and $a(t)$ is the scale factor. The overhead dot has been used to indicate the derivative with respect to cosmic time $t$. For the metric (6), the Friedmann equations can be expressed as [76]:
$H^{2}=\frac{1}{6f_{Q}}\left(\rho+\frac{f}{2}\right),\quad\dot{H}+3H^{2}+\frac{H\dot{f}_{Q}}{f_{Q}}=\frac{1}{2f_{Q}}\left(\frac{f}{2}-p\right).$ (7)
If $f(Q)$ is assumed to be $f(Q)=Q$ [76], one may obtain the standard field equations of the General Relativity. By considering $f(Q)=Q+F(Q)$, the above field equations (7) may also be expressed as
$3H^{2}=\rho+\frac{F}{2}-QF_{Q},\quad\dot{H}(2QF_{QQ}+F_{Q}+1)+\frac{1}{4}\left(2QF_{Q}-F+Q\right)=-\frac{p}{2}.$ (8)
The Eq. (8) may be seen as modifications of the General Relativity equations, incorporating an additional term that arises from the non-metricity ($Q$) of space-time. This term may exhibit a characteristics similar to a fluid component associated with dark sector of the fluid component of universe, i.e. $p_{Q}=p_{d}$ and $\rho_{Q}=\rho_{d}$. Consequently, using Eqs. (8), we may write
$3H^{2}=(\rho+\rho_{d}),\quad 2\dot{H}+3H^{2}=-(p+p_{d}).$ (9)
The pressure and density contributions to dark energy may be represented by $p_{d}$ and $\rho_{d}$ respectively, which may be caused by the non-metricity of spacetime. We define
${}\rho_{d}=\frac{F}{2}-QF_{Q},\quad p_{d}=2\dot{H}(2QF_{QQ}+F_{Q})-\rho_{d}.$ (10)
These equations may be used to obtain the solution of the system and thus the resulting Hubble parameter may be used for the identification of the cosmic evolution in the model. We proceed to set up the unified scenario of universe evolution in this $f(Q)$ model, where the transiting universe evolution may be described by the non-metricity scalar.
3 The cosmological solutions with the dark fluid in $f(Q)$ model
We proceed by taking the fluid which may represent the hydro-dynamically stable dark energy. The fluid has been assumed to have form [60, 65, 75]
$p_{d}=\alpha\rho_{d}-\rho_{0},$ (11)
where $\alpha$ and $\rho_{0}$ are some constants. For $\rho_{0}=0$, one may have the usual EoS $p_{d}=\omega\rho_{d}$ for $\alpha=\omega$. The parameter $\alpha$ may be visualized as constant speed of sound and the classical stability criterion constrains $\alpha$ as $0\leq\alpha\leq 1$. The causality and stability issues are important as far as the physical acceptability of a cosmological model is concerned [77, 78]. In a given medium, the sound speed describes the propagation speed of the acoustic (pressure) waves. The behavior of squared sound speed may be described by ${c_{s}}^{2}=\frac{\partial{p_{d}}}{\partial{\rho_{d}}}$. The uncontrolled growth of energy density perturbations may be prevented by ${c_{s}}^{2}>0$. The prevention of causality violation would lead to ${c_{s}}^{2}<1$.
In the subsequent sections, we study two different cases arising from Eq. (9). In general, the dark fluid may give rise to the transiting universe evolution (from matter dominated era into dark energy dominated era). We probe whether the dark fluid satisfying affine EoS yields this kind of scenario or not? For the affirmative case, we search for the functional form of $f(Q)$ in the $f(Q)$ gravity. In order to probe these issues, we assume $\rho=0$ in Eq. (9). As a consequence, the evolution of the Hubble parameter will be governed by the dark fluid (11) only. We nomenclate this considered scenario as Model I. In another case, we assume that the dark fluid (with energy density $\rho_{d}$ and pressure $p_{d}$) will account for the dark energy only and then assume that $\rho$ appears in the Eq. (9) is the matter energy density, that is $\rho=\rho_{m}$. As a consequence, $p=p_{m}=0$. These assumptions are motivated by the fact that matter (cold dark matter and baryonic matter) and dark energy are the dominant components of the present day universe. In the Model II, we probe for the universe’s evolution in $f(Q)$ gravity when the matter (following $p_{m}=0$) and dark energy (following Eq. (11)) are the dominant components of the universe in the model. We further describe the corresponding differences in universe evolution for these two cases. In other words, these two cases enable us to check the impact of the $\rho_{m}$ term in the cosmic dynamics yielded in these different models. The energy conservation equation for the dark fluid may be written as
$\dot{\rho}_{d}+3H(\rho_{d}+p_{d})=0.$ (12)
One may observe that for $\rho_{d}<\frac{\rho_{0}}{1+\alpha}$, one may have the phantom evolution of dark energy in model. For $\rho_{d}=\frac{\rho_{0}}{1+\alpha}$, the dynamics of constant vacuum energy may be visualized. For $\rho_{d}>\frac{\rho_{0}}{1+\alpha}$, the null energy condition will be satisfied in model, as a consequence the fluid distribution may either have the quintessence domination or the normal matter (composed of matter or radiation) domination. Based on these possible evolutionary dynamics, we probe the universe’s evolution under influence of dark fluid (given by Eq. (11)) in $f(Q)$ model.
3.1 Model I
In this model, we assume that the dark fluid (Eq. (11)) encompassed within $f(Q)$ gravity is the dominant component of the universe. For this case, $\rho=0$ in Eq. (9) leading to $\rho_{d}=3H^{2}$. From Eq. (11), it would lead to $F(Q)=-Q+C\sqrt{Q}$, where $C$ is an integration constant. The form of $f(Q)$ function may be written as
$f(Q)=C\sqrt{Q}.$ (13)
The above functional form of $f(Q)$ may describe the universe evolution yielded by the dark fluid (11). In other words, the affine EoS cosmology may belong to the class of cosmologies yielded by $f(Q)=C\sqrt{Q}$ gravity. The dark fluid conservation equation (12) may lead to the $\rho_{d}$ form depending on the scale factor as
${}\rho_{d}=\frac{\rho_{0}+(1+z)^{3(\alpha+1)}}{\alpha+1}.$ (14)
In this case, the Eqs. (9) and (14) would yield the Hubble parameter as
${}H(z)=H_{0}\sqrt{\frac{(1+z)^{3(1+\alpha)}+\rho_{0}}{\rho_{0}+1}},\quad\text{where,}\quad H_{0}=\sqrt{\frac{\rho_{0}+1}{3(1+\alpha)}}.$ (15)
The above Hubble parameter may describe the cosmic evolution in the $f(Q)$ gravity yielded by the fluid following affine equation of state (11). The observational constraints on the model parameters $\alpha,\rho_{0}$ and $H_{0}$ may describe the observational compatibility of the model along with the validity of causality principle in this model.
3.2 Model II
In this case, we take the universe composed of matter (cold dark matter and baryonic matter) with the dark energy governed by the dark fluid (11) in the $f(Q)$ gravity. For the present case, we take $\rho=\rho_{m}$ in Eq. (9), where $\rho_{m}$ denotes the energy density of matter satisfying EoS $p_{m}=0$. The conservation equation of matter $\dot{\rho}_{m}+3H\rho_{m}=0$ would lead to $\rho_{m}=\rho_{m0}a^{-3}$, where $\rho_{m0}$ is energy density of matter at $z=0$. And, $\rho_{d}$ is the dark energy density following Eq. (12). In this case, the Hubble parameter may be obtained as
${}H(z)=H_{0}\sqrt{\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda 0}+\Omega_{a0}(1+z)^{3(1+\alpha)}}$ (16)
where $\Omega_{m0}\equiv\frac{\rho_{m0}}{3H_{0}^{2}}$, $\Omega_{\Lambda 0}\equiv\frac{\rho_{0}}{3(1+\alpha)H_{0}^{2}}$ and $\Omega_{a0}\equiv\frac{\rho_{\alpha 0}}{3(1+\alpha)H_{0}^{2}}$. In the above expression, we use $\frac{a_{0}}{a}=1+z$ with $a_{0}=1$ for present day universe. At $z=0$ in present era, $H=H_{0}$ with constraint $\Omega_{m0}+\Omega_{\Lambda 0}+\Omega_{a0}=1$. This Hubble parameter may describe the cosmic dynamics in $f(Q)$ gravity where the fluids are composed of matter and dark energy satisfying $p_{m}=0$ and $p_{d}=\alpha\rho_{d}-\rho_{0}$ respectively. In this study, we ignore the contribution of radiation since it is observationally very small in the late-time universe [5]. We probe the observational compatibility of these models with the observational data and derive the constraint on model parameters.
4 Observational constraints and results
In this section, we focus on describing the data sets and methodology that are utilized to analyze the models. We also explain constraints on Model I (15) and Model II (16)’s model parameters. In other words, the consistency of the Hubble parameter (15) and (16) with the data sets given by the Cosmic Chronometer (CC) and combined data (consisting of Pantheon and CC) are used in the MCMC analysis. We constrain the parameters using $\chi^{2}$ minimization method and the MCMC method is implemented with the emcee Python library [79].
4.1 The Cosmic chronometer data
The model parameters are constrained using the Cosmic Chronometer (CC) data consisting of $31$ data points in the red-shift range of $0.07\leq z\leq 1.965$. These data are collected using the differential ages of galaxies approach [80, 81]. The cosmic chronometer observations are based on the principle introduced by Jimenez and Loeb [82]. It describes a relationship between red-shift $(z)$, cosmic time $t$, and the Hubble parameter $H(z)$ as $H(z)=\frac{-1}{(1+z)}\frac{dz}{dt}$. By minimizing the $\chi^{2}$ function, we determine the model parameters’ central values with $1\sigma$ errors in the emcee [79] package. The $\chi^{2}$ function has been taken as [75, 83, 84, 85]
${}\chi^{2}_{CC}(\theta)=\sum_{i=1}^{31}\frac{[H_{th}(\theta,z_{i})-H_{obs}(z_{i})]^{2}}{\sigma^{2}_{i}}.$ (17)
[…]
Figure 1: The best fit Hubble parameter curves with $z$ as compared to the $\Lambda CDM$ model.
It may be observed that the cosmic dynamics of the Model I and II are deviating from the $\Lambda$CDM model in the past where $z>1$. The transition redshift however remains close to the $\Lambda$CDM model. It is worthwhile to observe from these results that the Model I may describe the transiting universe evolution from the decelerated expansion into the accelerated expansion and even with the consideration of pressure-less matter in Model II, the cosmic dynamics remains similar as that of the Model I.
4.2 The Pantheon data
The Pantheon data is composed of supernovae type Ia (SNIa) observations identified in the low and high red-shift ranges [86]. This dataset includes $1048$ data points within the redshift interval $0.01\leq z\leq 2.26$. For the purpose of examining cosmic expansion of the observable universe, SNIa are widely adopted as the standard candles with their theoretical apparent magnitude $m_{th}(z)$ are given by:
${}m_{th}(z)=M+5\log_{10}\left[\frac{d_{L}(z)}{\text{Mpc}}\right]+25,$ (18)
where, $M$ is absolute magnitude, $c$ denotes the speed of light, $z$ represents the SNIa red-shift in the CMB rest frame and $d_{L}(z)=c(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ is the luminosity distance. The luminosity distance $d_{L}(z)$ is typically replaced with its Hubble-free counter part $(D_{L}(z)\equiv H_{0}d_{L}(z)/c)$. We can also rewrite equation (18) in the following way:
${}m_{th}(z)=M+5\log_{10}\left[D_{L}(z)\right]+5\log_{10}\left[\frac{c/H_{0}}{\text{Mpc}}\right]+25.$ (19)
In the $\Lambda$CDM model framework, a degeneracy will exist between $H_{0}$ and $M$ [87, 88]. We take $\mathcal{M}$ as a combination of these parameters [88]:
${}\mathcal{M}\equiv M+5\log_{10}\left[\frac{c/H_{0}}{\text{Mpc}}\right]+25$ (20)
where $H_{0}=h\times 100\ Km/s/Mpc$ and thus one can write $\mathcal{M}=42.39+M-5\log_{10}(h)$. For the Pantheon data, this parameter is used along with the relevant $\chi^{2}$ in the MCMC analysis defined as [88, 84, 83, 75, 85]
${}\chi^{2}_{P}=\nabla m_{i}C^{-1}_{ij}\nabla m_{j},$ (21)
where, $\nabla m_{i}=m_{obs}(z_{i})-m_{th}(z_{i})$, the theoretical apparent magnitude $m_{th}(z_{i})$ is calculated using equation (19) and the symbol $m_{obs}(z_{i})$ is used to denote the apparent magnitude of SNeIa observations at red-shift $z_{i}$ and $C_{ij}^{-1}$ is covariance matrix’s inverse. The luminosity distance is influenced by the Hubble parameter. Following this approach, we use equations (15) and (16) and the emcee package [79] to determine the maximum likelihood estimate using the joint (CC+Pantheon) dataset. For computing the maximum likelihood estimate using the joint data, we define the joint $\chi^{2}$ as $\chi^{2}_{CC}+\chi^{2}_{P}$.
[…]
CC $68.6^{+1.0}_{-1.0}$ $0.032^{+0.054}_{-0.054}$ $2.52^{+0.51}_{-0.70}$0.560 0.638 $13.51^{+0.08}_{-0.37}$ $1.145$ $-0.7068$
joint $68.9^{+1.8}_{-1.8}$ $0.056^{+0.11}_{-0.089}$ $2.78^{+0.67}_{-0.81}$ $23.806^{+0.011}_{-0.012}$0.580 0.65 $13.41^{+0.65}_{-1.04}$ $1.255$ $-0.7206$
Table 1: Model I: The constrained values with $q_{0},z_{t}$, $t_{0}$ $j_{0}$, where $H_{0}$ is in $Km/s/Mpc$.
[…]
CC $71.0^{+5.5}_{-8.2}$ $0.305^{+0.067}_{-0.056}$ $-1.33^{+0.73}_{-0.49}$ $-0.881$ $0.62$ $2.352$ $13.76$ $-0.921$
joint $68.7^{+1.9}_{-1.9}$ $0.322^{+0.051}_{-0.042}$ $-1.08^{+0.17}_{-0.13}$ $23.806\pm 0.014$ $-0.597$ $0.63$ $1.259$ $13.59$ $-0.7314$
Table 2: Model II: The constrained values with $q_{0},z_{t}$, $t_{0}$ and $j_{0}$, where $H_{0}$ is in $Km/s/Mpc$.
Figure 2: Model I: Marginalized $1$D and $2$D posterior contour map with median values of parameters using Joint data set.
Figure 3: Model II: Marginalized $1$D and $2$D posterior contour map with median values of parameters using Joint data set.
5 Cosmic dynamics and the physical behavior
In this section, we investigate the dynamical characteristics of the universe evolution in model I and II by using the observationally constrained values of model parameters.
5.1 Cosmographic parameter
The study of cosmography was initially introduced by Weinberg [89], who used a Taylor series to introduce the scale factor that rose around present time $t_{0}$. A detailed overview of cosmography may be reviewed for different aspects of cosmographic parameters [23]. The Hubble parameter ($H$), has been considered as a variable observable quantity. The behavior of the Hubble parameter (Fig. 1) highlights the expansion rate of the universe in considered models. The evolution of deceleration parameter $q$ is characterized through the second-order derivative of scale factor ($a$) [90]. To understand the universe’s cosmic evolution, the snap ($s$) and jerk ($j$) parameters serve as important tools. One of the cosmological parameters is the deceleration parameter that helps to characterize the universe’s expansion behavior. It is basically used to track down the rate of slowing down of the universe’s expansion. Mathematically, using the Hubble parameter and its derivative, it may be expressed as
${}q=-1-\frac{\dot{H}}{H^{2}}.$ (22)
This parameter serves as a tool to assess whether the universe is expanding at an increasing or decreasing rate. Depending on the value of $q$, different phases of cosmic expansion may be visualized. If $q$ is greater than zero $(q>0)$, the universe is in a decelerated phase. For $q<0$, the universe will expand at an accelerating rate. For $-1<q<0$ the expansion follows a power-law acceleration. A super-exponential expansion era occurs in the universe for $q<-1$, while a de Sitter expansion when $q=-1$ [91, 92, 93, 94]. For matter and radiation-dominated universe expansion, the deceleration parameter will be $q=\frac{1}{2}$ and $q=1$ respectively. By using (15), (16) and (22), we may obtain for Model I and II as
$q(z)=-1+\left({\frac{3(1+\alpha)(1+z)^{3+3\alpha}}{2(\rho_{0}+(1+z)^{3+3\alpha})}}\right),\ \ \text{and}$ (23)
$q(z)=-1+\frac{3(z+1)^{3}((\alpha+1)\Omega_{a0}(z+1)^{3\alpha}+\Omega_{m0})}{2((z+1)^{3}(\Omega_{a0}(z+1)^{3\alpha}+\Omega_{m0})+\Omega_{\Lambda 0})},$ (24)
respectively.
Figure 4: $q(z)$ versus $z$.
[…]
The deceleration parameter of these models provides insight into the universe’s transition from deceleration into acceleration of expansion. In model I, the present day value of deceleration parameter is $q_{0}=-0.56$ and $q_{0}=-0.58$ for CC and joint data respectively, and the value of transit red-shift are $z_{t}=0.638$ (for CC data) and $z_{t}=0.65$ (for joint data). The red-shift value at which $q=0$ is termed as the transit red-shift, since the behavior of $q$ is changing from positive to negative in the cosmic evolution. The accelerated universe expansion may be visualized in the model I for red-shifts $z<0.638$. For Model II, the transition from decelerated to accelerated expansion occurs at $z_{t}=0.62$ based on CC estimates and at $z_{t}=0.63$ using joint estimates. For CC estimates, the deceleration parameter’s current value is $q_{0}=-0.881$, while for joint estimates, it is $q_{0}=-0.597$. These negative values of deceleration parameters of model I and II are confirming that the universe is expanding with acceleration in the present era $(z=0)$. And, thus it has been observed that the non-metricity scalar may yield the transiting universe evolution in model I governed by $f(Q)=C\sqrt{Q}$. In the Model II, the dark energy introduced by the non-metricity may yield the accelerated expansion at late-times while the cold dark matter yields the decelerated universe expansion at early times in model.
The behavior of $q$ in Fig. (4) suggests that the model I will possess the matter-dominated era of universe expansion which may be confirmed with $(q\rightarrow\frac{1}{2})$ during early times. An accelerated phase $(q<0)$ in the present era and ultimately converges to the de Sitter scenario $(q=-1)$ as $z\rightarrow-1$. This model may explain the accelerated expansion universe at the present times, which is consistent with the observational data. It has been observed that the deceleration parameter of Model II is consistent with the current phase of accelerated expansion. In Fig. (4), the accelerated universe expansion of present times may reach to super-exponential expansion having $q<-1$ in the future. It is an important implication of the Model II, which is classically unstable due to $\alpha<0$.
The jerk parameter $(j)$ indicates the rate at which the universe’s acceleration or deceleration changes over time. The snap parameter $(s)$ calculates the rate at which the jerk parameter changes. These parameters are defined as follows:
${}\mathit{j}=\frac{1}{aH^{3}}\left(\frac{d^{3}a}{dt^{3}}\right),\ \ \mathit{s}=\frac{1}{aH^{4}}\left(\frac{d^{4}a}{dt^{4}}\right)$ (25)
In the present analysis, for $j$ and $s$, equation (25) can be rewritten in terms of red-shift as [95].
${}j(z)=q(z)\left(1+2q(z)\right)+\frac{dq}{dz}(1+z),\ \ \ \ s(z)=-\frac{dj}{dz}(1+z)-j(z)\left(2+3q(z)\right).$ (26)
For Model I, based on the constrained values from the CC data set, the snap parameter is $s_{0}=-2.12$ and the jerk parameter $j_{0}=1.14$. However, from the joint estimates, the jerk parameter value is $j_{0}=1.25$ and snap parameter value is $s_{0}=-2.03$. Similarly, for Model II, the value of the jerk parameter is $j_{0}=2.35$ (for CC data set) and $j_{0}=1.26$ (for joint data set). In the Tables (1) and (2), we summarize the present day values of the cosmological parameters. For Model I, the value of jerk parameter exhibits a decreasing trend from early to late times, eventually approaching to $1$. This behavior suggests that these models differ from $\Lambda$CDM in the early universe but the Model I would become similar to $\Lambda$CDM at later times.
5.2 The physical behavior of models
We highlight the physical behavior exhibited by the quantities such as the energy density and pressure. For the model parameters’ constrained values, the energy density has remained positive throughout the expansion history while the pressure may have undergone a recent transition to negative values. We calculate the energy density and pressure by utilizing equations (9), (15) and (16). In model I, the dark fluid having energy density $\rho_{d}$ will act as the total energy density of universe in model. And, the Model II possess the matter and the dark energy (due to dark fluid) will have the total energy density $\rho_{e}$.
The evolution of energy densities are depicted in Figs. (6) for Model I and II. The energy densities in these models will be increasing with red-shift $(z)$ (decreases with cosmic time $(t)$), remaining positive throughout expansion for both the models. In other words, the energy density will be decreasing from past to future in Model I. And, thus the energy density will have positive values and it rule out the existence of finite-time future singularities. In Model II, the energy density possess phantom nature. During the early universe, at large red-shift, the pressure commences with the positive values. In the later epoch along with the present epoch, the pressure becomes negative. It is observed from the graphical analysis that the non-metricity scalar will be responsible for the negative pressure necessary for the accelerated universe expansion in this late epoch.
Figure 5: Total energy density versus $\mathit{z}$
Figure 6: $\omega_{e}$ versus $\mathit{z}$
The EoS parameter may offer insights on the dominating fluid during the cosmic expansion of the universe and thus helps to determine the dynamics of the Universe. In the cosmological models, it may be commonly used to identify the nature of dark energy. The EoS parameter can be defined using the relationship between pressure $(p)$ and energy density $(\rho)$. It is mathematically expressed as $(\omega=\frac{p}{\rho})$. The cosmic characteristics, including the cold dark matter or dust dominated phase will have $\omega=0$ and the radiation-dominated period will have $\omega=\frac{1}{3}$. The transition into an accelerated cosmic era happens for the equation of state parameter satisfying $\omega<-\frac{1}{3}$. The cosmic acceleration period may have three regimes: the quintessence-dominated regime for $-1<\omega<-\frac{1}{3}$, the cosmological constant-like state for $\omega=-1$ and phantom-dominated regime $\omega<-1$. In the Fig. (6), we provide the graphical representation of the EoS parameter evolution in the considered models. In the Model I, the EoS parameter (defined as $\omega_{e}=p_{d}/\rho_{d}$) values are estimated as $\omega_{e}=-0.706$ and $\omega_{e}=-0.720$ for CC and joint dataset respectively at $z=0$. The Model II possess matter and dark energy, as a consequence, we define $\omega_{e}=p_{e}/\rho_{e}$, where the total energy density and pressure are denoted by $\rho_{e}$ and $p_{e}$ respectively. In the Model II, the effective EoS parameter values are obtained $\omega_{e}=-0.921$ (for CC dataset) and $\omega_{e}=-0.73$ (for joint dataset) at $z=0$. The Fig. (6) is revealing that the effective EoS parameter may correspond to the quintessence kind of dark energy in these models at present times. In future, as $z\rightarrow-1$, the cosmological constant-like behavior may be visualized in the Model I. On the other hand, in the Model II, the phantom kind of dark energy may evolve during the cosmological evolution. In the present era, the dark energy is a dynamical one which is varying with time in these models.
These models possess the past of matter-dominated expansion which are confirmed by Fig. (6) having $\omega=0$ during early times. The cosmic evolution is under the domination of quintessence kind of dark energy having $-1<\omega<-\frac{1}{3}$ in the present epoch. However, the future evolution in both of these models may be different. For these evolution scenarios, we observe the role of non-metricity scalar having geometrical origin in the $f(Q)$ model. In the model I, the non-metricity is in charge of the matter dominated expansion and the dark energy dominated expansion. In this sense, the unified scenario has been confirmed in the affine EoS induced model of $f(Q)$ gravity. On the other hand, when the cold dark matter has been considered with the dark fluid satisfying affine EoS, the $f(Q)$ gravity model yields the identical matter evolution as of the $\Lambda$CDM model but the dark energy behavior will be clearly different. It is an important difference between the cosmic dynamics possessed by models without and with cold dark matter. In the model without dark matter, the fluid with affine EoS unifies the decelerated and accelerated expansion. While in the model with cold dark matter and fluid having affine EoS, the matter dominated era may be the same but the phantom evolution of dark energy may not be ruled out in $f(Q)$ gravity.
5.3 Age of the Universe
The age of the universe is an important criterion to test the viability of the model as compared to the observable universe. The cosmic age $t(z)$ may be expressed in terms of the red-shift $z$ as  [96]
${}t(z)=\int_{z}^{\infty}\frac{dz^{\prime}}{(1+z^{\prime})H(z^{\prime})}.$ (27)
The present age $(t_{0})$ at $(z=0)$ has been calculated numerically using the integral consisting of the Hubble parameters $H(z)$ provided in equations (15) and (16).
In the Model I, the universe’s current age is $t_{0}$ (at $z=0)=13.51^{+0.08}_{-0.37}$ Gyr for CC data sets and $t_{0}=13.41^{+0.65}_{-1.04}$ Gyr for joint data sets, both of which are close with the current age values of the $\Lambda$CDM model (having $t_{0}=13.79$ Gyr) obtained from Planck results [5].
In model II, the universe present age value is $t_{0}=13.76$ Gyr and $t_{0}=13.59$ Gyr for CC and joint estimates respectively, both of which are again close with the current age values of the $\Lambda$CDM model mentioned above.
6 Conclusions
We probe the possibility to describe the unified scenario of universe expansion in $f(Q)$ gravity. The late-time accelerated expansion of the Universe has been investigated in the $f(Q)$ gravity framework with an affine equation of state (EoS) $(p_{d}=\alpha\rho_{d}-\rho_{0})$. In the $f(Q)$ model with affine EoS, we show that the unified scenario may be realized where the non-metricity may yield the transition from the decelerated expansion into accelerated expansion. In this model, a particular functional form $f(Q)=C\sqrt{Q}$ of $f(Q)$ gravity has been derived and it is shown that the matter dominated expansion may also be explained with the non-metricity scalar in the model. In other words, we show that the unified cosmological scenario yielded by the affine EoS fluid may belong to class of $f(Q)=C\sqrt{Q}$ cosmologies in the $f(Q)$ gravity framework. Further, we also study the model composed of cold dark matter and the dark fluid satisfying the affine EoS. Descriptive differences between these models have been studied.
We use the observational data of the CC and supernovae type Ia Pantheon data for the MCMC analysis based on the Hubble parameters (15) and (16) of the derived models. A summary of the constrained values extracted from the MCMC application in the models are provided in Tables (1) and (2).
The evolution dynamics in the present study has been examined by study of the cosmological parameters. The behavior of the deceleration parameter has been investigated. The evolution and dominant cosmic components during various epochs are reflected in the deceleration parameter behaviors of both models. The universe’s phase recently transitioned from decelerated to accelerated as indicated by the deceleration parameter’s evolution curve. In Model I, the estimated present deceleration parameter values are $q_{0}=-0.56$ and $q_{0}=-0.58$ for the CC and joint analyses respectively. For Model II, we obtain the deceleration parameter values at present as $q_{0}=-0.881$ and $q_{0}=-0.597$ based on the CC and joint data respectively. A negative deceleration parameter at $z=0$ demonstrates that cosmic expansion is accelerating for both models. At late-times, the negative pressure introduced by the non-metricity may be responsible for the accelerated universe expansion.
In the considered models, the physical behavior has been elaborated on the basis of energy density, pressure and EoS parameters for the model parameters’ constrained values. The energy density remains positive whereas pressure exhibits a change from positive to negative values. It is observed in the models that the cosmic accelerated expansion is caused by this negative pressure available in the late era of the universe expansion. The universe in Model I may approach to the $\Lambda$CDM limits in future as $z\rightarrow-1$. The varying dark energy evolution may correspond to the cosmological constant-like behavior in the asymptotic era having $z\rightarrow-1$. While the energy density of the universe in Model II will increase in future due to phantom-like characteristics. The parameter constraints subjected to the CC and Pantheon data suggest that Model I is classically stable while Model II is classically unstable. It is an important distinction as far as physical acceptability of these models are concerned.
In particular, for the model I, at $z=0$, $\omega_{e}=-0.706$ and $\omega_{e}=-0.720$ for CC and joint data respectively. These values at the present time suggest the existence of a quintessence-type dark energy in the model. Similarly, for the Model II, the EoS parameter values are obtained $\omega_{e}=-0.921$ (for CC dataset) and $\omega_{e}=-0.731$ (for joint dataset) at $z=0$. The graphical behavior reveals that the EoS parameter for dark energy exhibits quintessence kind of dark energy at present times. In summary, the $f(Q)$ model with affine EoS is consistent with the observations of cosmic acceleration. The violation of null energy condition in the Model II may not prohibit from the existence of finite-time future singularity.
For the $\Lambda$CDM model, the cosmological parameters are $q=-1$, $\omega=-1$, $j=1$, Snap $(s)=1$ during the asymptotic era as $z\rightarrow-1$. In other words, in the de-Sitter model of cosmic expansion, where $H$ remains constant and $q=-1$, $j=1$ and $s=1$. This indicates that the universe is undergoing exponential acceleration with the rate of change of both acceleration and jerk increasing. In this context, for the classically stable Model I (Eq. 15), the dark energy equation of state (EoS) parameter ($\omega_{e})$ at present is near $-0.7$ and the deceleration parameter is approximately $-0.56$ (for CC data) and $-0.58$ (for joint data). In this scenario, the universe’s acceleration is attributed to dynamical dark energy and the jerk parameter is close to $\Lambda$CDM limit $(j=1)$ while the snap parameter is negative for the median values of model parameters. The Tables (1) and (2) highlight a summary of the results on cosmographic parameters. It is worthwhile to mention that these quantities will eventually align with the $\Lambda$CDM limits in the distant future. However, at present times, the cosmic dynamics of these models are differing from the $\Lambda$CDM model.
The age of the universe is also estimated in the context of present $f(Q)$ gravity models. For model I, the value is $t_{0}=13.51$ Gyr and $t_{0}=13.41$ Gyr for CC and joint estimates respectively. For model II, the universe’s present age value is $t_{0}=13.76$ Gyr (for CC data) and $t_{0}=13.59$ Gyr (for joint data).
In summary, we show that the observationally compatible unified scenario may be realized in the $f(Q)$ gravity framework with the affine EoS. The non-metricity scalar may explain the matter dominated era of the past along with the dark energy dominated era of the observable universe. A descriptive difference between two classes of models have illustrated that the present day matter density of the considered models are approximately the same but the Model I is classically stable while Model II is classically unstable. The dynamical dark energy in Model I is of quintessence kind with $\Lambda$CDM limit in future while the dynamical dark energy of Model II may have the phantom nature in future.
Acknowledgments
GPS and AS are thankful to the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India for support under Visiting Associateship program.